OpenStudy (anonymous):

how do i do this problem from factored form 4x^2+28x+40

OpenStudy (anonymous):

FACTOR OUT THE FOUR

OpenStudy (anonymous):

4(x^2+5)(x+2)

OpenStudy (anonymous):

(2*x)^2+2*2*x*7+(7^2)+40-(7^2) Using formula (A+B)^2=A^2+2*A*B+B^2 [(2*x+7)^2]+40-49 [(2*x+7)^2]-9 [(2*x+7)^2]-(3)^2 using formula (a)^2-(b)^2=(a+b)*(a-b) We get [2*x+7+3}*[2*x+7-3] [2*x+10]*[2*x+4]